OpenStudy (ahsome):

A man is hauling a box of mass 100kg up a 35 degree incline by a rope attatched to the top of the box. If the rope makes an angle of 20 degrees to the incline, and the coefficient of friction between the box and the incline is 0.65, calculate the force applied by the man to keep the box moving at a constant speed.

OpenStudy (ahsome):



OpenStudy (shamim):

Weight of box= mg=100*9.8N=?

OpenStudy (ahsome):

\[980N\]

OpenStudy (shamim):

This weight will b acting just downward. Right?

OpenStudy (ahsome):

Yes

OpenStudy (ahsome):

Should I type what I did?

OpenStudy (shamim):

This weight 980 N has 2 perpendicular component

OpenStudy (ahsome):

Yes. One perpendicular to the plane, and one parallel to the plane (right)?

OpenStudy (shamim):

Okb. U can type wt u did. Plz!!!

OpenStudy (ahsome):

:)

OpenStudy (shamim):

Ya. Right

OpenStudy (ahsome):

Ok, first i found what gravity was: \[Fg=m\times g\]\[Fg=100*9.8\]\[Fg=980N\]

OpenStudy (shamim):

Ok

OpenStudy (ahsome):

I then find the force pushing down, parallel to the plane \[Fp=Fg\times sin\theta\]\[Fp=980N*sin(35)\]\[Fp\approx561.9N\]

OpenStudy (ahsome):

I then split the force of the pull into its \(x\) and \(y\) components

OpenStudy (ahsome):

\[Fa_x=Fa\times \sin(20)\]\[Fa_y=Fa*\cos(20)\]

OpenStudy (shamim):

No

OpenStudy (ahsome):

Oh, K. What do I do then

OpenStudy (anonymous):

Again, to get frictional force, you need to find the N, which is mgsin@ or mgsin35, that is 100*9.8*cos35=802.8N. Frictional force or R is $N or 0.65*802.8=521.8N, this is the frictional force due to the slopes friction. Next, take account of the driving force, which is the one acting at 20degrees, so the Fd=Tension*cos20 or Tcos20 . Finally, we can now use the Newtons formula Fd -R =ma. Given that the speed is constant, then a= 0, rewriting the formula as Fd=R Take note that the forces acting downward apart from the friction one is also that of weight, which is mgsin35 or 100*9.8*sin35=562.1N Then, we shall have Tcos20= 562.1 + 521.8. And T=1153N.

OpenStudy (shamim):

U split pulling force along the plane nd perpendicular to the plane

OpenStudy (ahsome):

@shamim, isn't that what I did? After I said I was gonna split it (I never splt Fp)

OpenStudy (anonymous):

It is just that the tension acts parallel to the plane, when tension is multiplied by cos20, which is what I made as driving force.

OpenStudy (shamim):

U r right @Ahsome

OpenStudy (ahsome):

Phew. Got to do something. Can I explain the rest in a bit?

OpenStudy (shamim):

Ok. Keep going

OpenStudy (shamim):

I think u hv 2 unequal but oposit forces

OpenStudy (shamim):

U can minus frm bigger

OpenStudy (anonymous):

Me?

OpenStudy (shamim):

Why u r silent @Ahsome

OpenStudy (ahsome):

@shamim, I said I needed to do something for a little bit. Can I come back and we can discuss in a little bit (dinner)

OpenStudy (shamim):

Ok

OpenStudy (anonymous):

But, didn't my explanation also facilitated a bit?

OpenStudy (anonymous):

Bon Appetite anyway. Over here is still 2:09pm.

OpenStudy (shamim):

Which country u r? @Hoslos

OpenStudy (ahsome):

Back You guys still here, @shamim and @Hoslos?

OpenStudy (shamim):

Ok. I m here

OpenStudy (ahsome):

Ok. So, we have the following things: \[Fg=980N\] \[Fa_y=Fa\times \sin(20)\]\[Fa_x=Fa\times \cos(20)\] \[Fp=561.N\]

OpenStudy (ahsome):

Now, last thing we need to do is find \(Fn\). Normally, \(Fn\) is equal to \(Fg\times \cos(\theta)\), but in this situation part of the pull force pushes the object up. Whenthe object goes up, the \(Fn\) becomes smaller. In the end: \[Fn=Fg\times \cos\theta - y\]\[Fn=980\times cos(35)-y\]

OpenStudy (ahsome):

Ok. So, we have the following things: \[Fg=980N\] \[Fa_y=Fa\times \sin(20)\]\[Fa_x=Fa\times \cos(20)\] \[Fp=561.N\] \[Fn=980\times cos(35)-y\]

OpenStudy (ahsome):

We know the equation that \(Fr=Fn\times µ\) Where Fr is friction, Fn is normaly and µ is co-efficient of friction

OpenStudy (anonymous):

I am from southern Africa. What about you @shamim?

OpenStudy (ahsome):

Ok. So, we have the following things: \[Fg=980N\] \[Fa_y=Fa\times \sin(20)\]\[Fa_x=Fa\times \cos(20)\] \[Fp=561.N\] \[Fn=980\times cos(35)-y\] \[Fr=Fn\times µ\]

OpenStudy (ahsome):

Welcome back @Hoslos :D

OpenStudy (ahsome):

Let me continue...

OpenStudy (anonymous):

Yes!

OpenStudy (ahsome):

Since we know that the object is moving at a constant velocty, then whatever is pushing up has to be equal to pushiing down. The force pushing up is the x value of the pulling force. The forces pushing down (the plane) are Friction and the part of gravity parallel to the plane

OpenStudy (ahsome):

\[Fr+Fp=x\]

OpenStudy (anonymous):

When you do Fgsin35 it is the force of weight acting parallel to the slope, which is correct.

OpenStudy (ahsome):

Yes

OpenStudy (anonymous):

Correct for forces acting down the plane.

OpenStudy (anonymous):

But the force that acts up the plane is Tcos20, because it is said to act at a given angle.

OpenStudy (ahsome):

We know that \[Fr=Fn\times µ\]. So lets put that into the equation: \[Fr+Fp=x\]\[Fn×µ+Fp=x\]

OpenStudy (anonymous):

Eventually Tcos20= friction + force of weght parallel to plane.

OpenStudy (ahsome):

Exactly

OpenStudy (ahsome):

\[Fn×µ+Fp=x\]

OpenStudy (shamim):

Ya

OpenStudy (ahsome):

Lets slowly sub everything we know

OpenStudy (shamim):

Right!!!

OpenStudy (anonymous):

So 561 is of weight and 803 is friction, right?

OpenStudy (ahsome):

We don't know friction yet (I haven't showed that yet ;) ) Put yes, 561 is the force parallel to the plane

OpenStudy (anonymous):

I mean friction is 522, right?

OpenStudy (ahsome):

I am not sure

OpenStudy (anonymous):

To get this, you first find force to the y-plane, being mgcos35 or 100*9.8*cos35=802.8, which is the contact force. This is the first step when going for friction. Are you ok up to this point?

OpenStudy (ahsome):

Yes, but you need to take away the y component of the pulling force, since that does part of the work for Fn, so Fn is smaller

OpenStudy (shamim):

Correct @Ahsome

OpenStudy (anonymous):

Hmm, I am seeing that we different ways of solving this problem. But I believe mine is simpler. The way I learnt was that when you have an inclined plane and want to find frictional force, you first find the y-plane force which is mgcos$ to get the normal contact force and after that we do frictional force=Mu*N , where Mu is the coefficient of fricion and N is the contact force.

OpenStudy (ahsome):

Exactly. That works if no force moves in the same direction as Fn

OpenStudy (ahsome):

But the thing is, the pulling force does move in that direction (the y component)

OpenStudy (ahsome):

So you need to take that into consideration aswell

OpenStudy (anonymous):

So, my point after finding the 802.8N, was to then base from the formula R=MuN or 802.8*0.65=521.8N - the frictional force.

OpenStudy (ahsome):

But remember, the normal force is smaller than 802.8

OpenStudy (ahsome):

Right @shamim? (I hope I am)

OpenStudy (shamim):

Right@Ahsome

OpenStudy (anonymous):

Yes, because you have to subtract with the Tsin20.

OpenStudy (ahsome):

Yup. But you don't know T?

OpenStudy (anonymous):

That is what thewuestion wants.

OpenStudy (ahsome):

Yup. So how could you assume that Fn is 801

OpenStudy (anonymous):

What I am thinking is that we make that normal contact force subtracting the Tsin20 as one equation and we multiply with the Mu then add with the forceof weight parallel to plane. This will be the equation of the forces acting downwards. And we equate to Tcos20, which is the force acting upwards, making as find T the subject of the formula, because we will have an equation.